OpenStudy (anonymous):

I need some to show me step by step on adding expressions example: (1/2x)+(1/2)= (4/x+3)

OpenStudy (anonymous):

i don't think this is an addition problem, i think it is and equation to solve' am i right?

OpenStudy (anonymous):

bring the x's and constants to one side so it would be 1/2x-4/x=3-1/2

OpenStudy (anonymous):

ok so now you need to find the common denominator. 1/2x-8/2x= -7/2x

OpenStudy (anonymous):

-7/2x= 5/2 you can cross multiply which will get you -14=10x divide both sides by 10 and you will get -14/10=x you can simplify it more to get -7/5=x